The following selected transactions relate to investment activities of Ornamental Insulation Corporation during 2021. The company buys debt securities, intending to profit from short-term differences in price and maintaining them in an active trading portfolio. Ornamental’s fiscal year ends on December 31. No investments were held by Ornamental on December 31, 2020.

Mar. 31 Acquired 8% Distribution Transformers Corporation bonds costing $510,000 at face value.
Sep. 1 Acquired $1,230,000 of American Instruments' 10% bonds at face value.
Sep. 30 Received semiannual interest payment on the Distribution Transformers bonds.
Oct. 2 Sold the Distribution Transformers bonds for $590,000.
Nov. 1 Purchased $1,950,000 of M&D Corporation 6% bonds at face value.
Dec. 31 Recorded any necessary adjusting entry(s) relating to the investments. The market prices of the investments are
American Instruments bonds $ 1,181,000
M&D Corporation bonds $ 2,021,000

(Hint: Interest must be accrued.)

Required:
1. Prepare the appropriate journal entry for each transaction or event during 2021, as well as any adjusting entries necessary at year end.
2. Indicate any amounts that Ornamental Insulation would report in its 2021 income statement, 2021 statement of comprehensive income, and 12/31/2021 balance sheet as a result of these investments. Include totals for net income, comprehensive income, and retained earnings as a result of these investments.

Required 1 Required 2 Indicate any amounts that Ornamental Insulation would report in its 2021 income statement, 2021 statemeI need help with the Cash under current assets.
Answer #1

Cash paid for acquiring 8% Distribution Transformer Corporation Bond at $510,000 at fair value so cash here is -$510,000.

On 30 Sept Received semiannual interest on bonds which will be 8% on 510,000 for 6 months which is $20,400

On 2 Oct sold 8% Distribution Tranformer Bond for $590,000. Here the cahs balance will be after selling = -510,000 + 20,400 + 590,000 = 100,400

Also on Sept 1 he acquired $1,230,000 American Investment 10% bond at face value. So again the cash balance will reduce. so now Cash balance = 100,400 - 1,230,000 = -$1,1129,600

Again on 1 Nov he purchased$1,950,000 of M&D Corporation 6% Bonds at face value.

So his cash balance will be = -1,129,600 - 1,950,000 = -$3,079,600

At the end of Dec 2020 he sold all the investment at the market price so he sold American Investment for $1,181,000 and M&D Corporation Bond for $2,021,000 which totals the incoming cash $3,202,000.

Cash Balance = - $3,079,600 + $3,202,000

= $122,400

So Cash BAlance will be $122,400